OpenStudy (anonymous):

Find the area of a square pizza if one side equals 11 centimeters. 0.44cm^2 √121cm^2 44cm^2 121cm^2

OpenStudy (anonymous):

121 cm^2

OpenStudy (anonymous):

\[area=s^2\] \[11^2\] 121

OpenStudy (angela210793):

A=x^2=11^2=121cm^2